6(b) Write the dual of the following LPP after reducing it to canonical form. (4) 

 Min 3 1 4 2 3 3 Z = x + x + x

 Subject to 

 2x1 + 4x2 =12

 11 5x1 + 3x3 ≥

 6x1 + x2 ≥ 8

 x1

, x2

, x3 ≥ 0


1
Expert's answer
2021-04-15T07:20:29-0400

Solution.

"z=3x_1+4x_2+3x_3 - min, \newline
2x_1+4x_2=12,\newline
5x_1+3x_3>=11, \newline
6x_1+x_2>=8, \newline
x_1,x_2,x_3>=0."

Reduce it to canonical form:

"2x_1+4x_2=12,\newline
-5x_1-3x_3+x_4=-11, \newline
-6x_1-x_2+x_5=-8. \newline"

Introduce new variables "x_4, x_5,x_6:"

"2x_1+4x_2+x_6=12,\newline
-5x_1-3x_3+x_4=-11, \newline
-6x_1-x_2+x_5=-8. \newline"

then "z = Mx_6=\newline =M(12-2x_1-4x_2)=\newline =-2Mx_1-4Mx_2+12M - min"

Introduce new variable x_0:

"x_0=-2x_1-4x_2."

We will have:

"x_0=12-2x_1-4x_2,"

"x_4=-11+5x_1+3x_3,\newline
x_5=-8+6x_1+x_2,\newline
x_6=12-2x_1-4x_2."

This plan is not optimal because there are negative elements in the expression for "x_0" . Select a new variable "x_1," then

"x_1=\frac{11}{5}-\frac{3}{5}x_3+\frac{1}{5}x_4."

From here

"x_0=\frac{38}{5}-4x_2+\frac{6}{5}x_3-\frac{2}{5}x_4,"

"x_5=\frac{26}{5}+x_2-\frac{18}{5}x_3+\frac{6}{5}x_4,\newline
x_6=\frac{38}{5}-4x_2+\frac{6}{5}x_3-\frac{2}{5}x_4."

The expression for "x_0" has negative elements, the plan is not optimal. The largest coefficient is near "x_2" , then choose "x_2" as a new variable.

"x_2=\frac{19}{10}+\frac{3}{10}x_3-\frac{1}{10}x_4-\frac{1}{4}x_6,
\newline
x_0=0+x_6,
\newline
x_1=\frac{11}{5}-\frac{3}{5}x_3+\frac{1}{5}x_4,\newline
x_5=\frac{71}{10}-\frac{33}{10}x_3+\frac{11}{10}x_4-\frac{1}{4}x_6."

The expression for "x_0"

has positive elements, the plan is optimal.

We will deduce variables "x_4, x_5,x_6" from basis:

"x_0=\frac{71}{5}+\frac{12}{5}x_3+\frac{1}{5}x_4,"

"x_1=\frac{11}{5}-\frac{3}{5}x_3+\frac{1}{5}x_4,"

"x_2=\frac{19}{10}+\frac{3}{10}x_3-\frac{1}{10}x_4,\newline
x_5=\frac{71}{10}-\frac{33}{10}x_3+\frac{11}{10}x_4."

Optimal plan:

"x_1=2\frac{1}{5}, x_2=1\frac{9}{10}, x_3=0."

So, min "z(x)=14\frac{1}{5}."

Answer. min "z(x)=14\frac{1}{5}."